one man is five times as old as his brother. fifteen years later he will be twice as old. how old are the two men now?
Answer by ankor@dixie-net.com(22740)

Let m = man's present age
Let b = brother's
:
one man is five times as old as his brother.
m = 5b
:
fifteen years later he will be twice as old.
m + 15 = 2(b+15)
m + 15 = 2b + 30
m = 2b + 30 - 15
m = 2b + 15
Replace m with 5b
5b = 2b + 15
5b - 2b = 15
3b = 15
b = 15/3
b = 5 yrs is brother's present age
then
5(5) = 25 yrs is man's
:
:
Check this in the statement:
" fifteen years later he will be twice as old."
25 + 15 = 2(5 + 15)
40 = 2(20)